# Node Perturbation Can Effectively Train Multi-Layer Neural Networks

### v6

@@ Appendix F Experimental details
Experiments were performed using fully-connected or convolutional neural networks with leaky ReLU activation functions and a categorical cross-entropy (CCE) loss on the one-hot encoding of class membership. The Adam optimizer with default parameters $\beta_{1}=0.9$, $\beta_{2}=0.999$, $\epsilon=1e-7$, was used for optimization. All experiments were run on an HP OMEN GT13-0695nd desktop computer with an NVIDIA RTX 3090 GPU. The experiments required about a week of total computation time and several more weeks of computation time during exploratory experiments.
Table 1 provides details of the employed neural network architectures.
Table 1: The neural network architectures consist of fully connected (FC) and convolutional (Conv) layers. All layers except the output layer are followed by a leaky ReLU transformation. Convolutional layer size is given as height×$\times$×width×$\times$×output channels, stride.
@@
Six hidden layers 6 $\times$ FC 1024
FC 10
Nine hidden layers 9 $\times$ FC 1024
FC 10
ConvNet Conv 3$\times$3$\times$16, 2
Conv 3$\times$3$\times$32, 2
Conv 3$\times$3$\times$64, 1
FC 1024
FC 10
ConvNet Tiny ImageNet Conv 5$\times$5$\times$32, 2
Conv 3$\times$3$\times$64, 2
Conv 5$\times$5$\times$128, 1
@@
Method 1 layer 3 hidden layers 6 hidden layers 9 hidden layers ConvNet ConvNet Tiny ImageNet
NP $1.0\times 10^{-3}$ $1.0\times 10^{-4}$
DNP $1.0\times 10^{-3}$ $1.0\times 10^{-3}$ $5.0\times 10^{-3}$ $5.0\times 10^{-3}$ $1.0\times 10^{-3}$
ANP $1.0\times 10^{-4}$ $1.0\times 10^{-4}$
DANP $1.0\times 10^{-3}$ $5.0\times 10^{-3}$ $5.0\times 10^{-3}$ $1.0\times 10^{-4}$ $1.0\times 10^{-3}$
INP $1.0\times 10^{-4}$ $1.0\times 10^{-4}$
DINP $1.0\times 10^{-3}$ $5.0\times 10^{-3}$ $5.0\times 10^{-3}$ $1.0\times 10^{-4}$ $1.0\times 10^{-3}$
BP $1.0\times 10^{-3}$ $1.0\times 10^{-4}$ $1.0\times 10^{-3}$ $1.0\times 10^{-4}$
DBP $1.0\times 10^{-3}$ $1.0\times 10^{-3}$ $1.0\times 10^{-3}$
Appendix G Alignment as a function of noisy forward passes
@@
Appendix H Impact of perturbation scale on performance
Appendix I Peak accuracies
Table LABEL:tab:peakacc shows peak accuracies for all algorithms used in the CIFAR-10 experiments and Table LABEL:tab:peakacc_tin shows peak accuracies for the Tiny ImageNet experiment.
@@ Appendix I Peak accuracies
1 layer 3 hidden layers 6 hidden layers 9 hidden layers ConvNet
Method Train Test Train Test Train Test Train Test Train Test
BP 41.3 39.3 59.6 51.5 100.0 62.1
DBP 44.9 40.9 99.6 54.3 100.0 60.5
NP 38.5 38.2 28.1 28.1 45.2 44.5
DNP 42.8 40.2 44.9 42.1 46.7 43.1 40.5 38.6 50.2 47.9
ANP 30.4 30.6 52.7 49.9
DANP 48.5 44.4 47.3 44.2 44.8 42.7 58.2 52.6
INP 32.0 32.0 52.3 49.9
DINP 52.6 46.3 53.0 46.6 52.0 45.3 57.9 52.2
@@
Appendix J SARCOS experiment
Figure 7 shows train and test loss for (D)NP, (D)ANP and (D)BP for the SARCOS dataset [Vijayakumar and Schaal, 2000]. The task is to solve inverse dynamics problem for a seven degrees-of-freedom SARCOS anthropomorphic robot arm by mapping from a 21-dimensional input space (7 joint positions, 7 joint velocities, 7 joint accelerations) to the corresponding 7 joint torques. The network is a fully connected architecture with two 32-unit hidden layers and was trained with MSE loss and the leaky ReLU activation function. The network was trained for 500 epochs and results were averaged over three random seeds. A learning rate search for each algorithm was started at $10^{-7}$, doubling the learning rate until the algorithm became unstable. A decorrelation learning rate of $10^{-4}$ was used for DNP and DANP and $10^{-3}$ for DBP.
Appendix K Deeper CIFAR architectures
This section describes the scaling of DNP, DANP and DINP to deeper fully connected networks as well as convolutional neural networks.
Figure 8 (Top panel) compares the performance of DNP, DANP and DINP for 3, 6 and 9 hidden layers. Performance for DINP and DANP scales better with network depth than performance for DNP, again highlighting the improvement these novel formulations provide.
Figure 8 (Bottom panel) shows that, when training a convolutional neural network on CIFAR-10, DANP and DINP outperform NP, though relative to BP, a large gap in train accuracy emerges. Also, the NP-style algorithms lag behind BP more than they did in the fully-connected networks. All decorrelated NP variants massively outperform their non-decorrelated counterparts in both train and test accuracy. Interestingly, (D)ANP outperforms (D)NP in this experiment as well.
Figure 8: Performance comparisons. Curves report mean train and test accuracy. Shaded areas indicate the maximal and minimal accuracy obtained for three random seeds. Top: Performance of DNP, DANP and DINP for 3-, 6- and 9-hidden-layer networks. Bottom: Performance of different node perturbation and backpropagation variants when training a convolutional neural network on CIFAR-10.

### v7

@@ Appendix D Experimental details
Experiments were performed using fully-connected or convolutional neural networks with leaky ReLU activation functions and a categorical cross-entropy (CCE) loss on the one-hot encoding of class membership. The Adam optimizer with default parameters $\beta_{1}=0.9$, $\beta_{2}=0.999$, $\epsilon=1e-7$, was used for optimization. All experiments were run on an HP OMEN GT13-0695nd desktop computer with an NVIDIA RTX 3090 GPU. The experiments required about a week of total computation time and several more weeks of computation time during exploratory experiments. Table 1 provides details of the employed neural network architectures.
[…]
@@
ConvNet Tiny ImageNet Conv $\times5$$\times$32, 2
Conv $\times3$$\times$64, 2
Conv $\times5$$\times$128, 1
@@
Method 1 layer 3 hidden layers ConvNet Tiny ImageNet
NP $\times 10^{-3} 1.0×10$$\times 10^{-4}$
DNP $\times 10^{-3} 1.0×10$$\times 10^{-3}$ $1.0\times 10^{-3}$
ANP $1.0\times 10^{-4}$
DANP $\times 10^{-3} 1.0×10$$\times 10^{-3}$
INP $1.0\times 10^{-4}$
DINP $\times 10^{-3} 1.0×10$$\times 10^{-3}$
BP $\times 10^{-3} 1.0×10$$\times 10^{-4}$ $1.0\times 10^{-4}$
DBP $\times 10^{-3} 1.0×10$$\times 10^{-3}$
Appendix E Alignment as a function of noisy forward passes
@@
Appendix F Peak accuracies
Table 3 shows peak accuracies for all algorithms used in the CIFAR-10 experiments and Table LABEL:tab:peakacc_tin shows peak accuracies for the Tiny ImageNet experiment.
@@ Appendix F Peak accuracies
1 layer 3 hidden layers
Method Train Test Train Test
BP 41.3 39.3 59.6 51.5
DBP 44.9 40.9 99.6 54.3
NP 38.5 38.2 28.1 28.1
DNP 42.8 40.2 44.9 42.1
ANP 30.4 30.6
DANP 48.5 44.4
INP 32.0 32.0
DINP 52.6 46.3
@@
Appendix G Computational overhead of decorrelation
This section describes the theoretical amount of extra computation and memory that is required when applying decorrelation two a fully connected layer, a convolutional kernel and the amount of measured extra memory use in a three-layer fully connected network. Tables 5 and 6 display FLOPs and memory requirements for a fully connected layer and a convolutional layer, respectively. Table 7 shows measured memory consumption for a three-layer fully connected network for the various algorithms.
Table 5: Approximate FLOPS required for a forward pass through a fully connected $1000\times 1000$ layer, batch size $1000$, for BP, DBP, and DANP
Algorithm Approx. FLOPS Brief Description of Calculation Approx. Memory Required
BP Total 6,000,000,000 forward pass + backward pass + weight update $6\times\text{batch\_size}\times\text{in\_dim}\times\text{out\_dim}$ $\mathcal{O}\!\big(\text{batch\_size}\times(\text{in\_dim}+\text{out\_dim})+\text{in\_dim}\times\text{out\_dim}\big)$
DBP Total 14,000,000,000 forward $R\times z +$ backward $R^{\top}\times e +$ weight update $XX^{\top} +$ R update $6\times\text{batch\_size}\times\text{in\_dim}\times\text{out\_dim}+4\times\text{batch\_size}\times\text{in\_dim}^{2}+2\times\text{in\_dim}^{3}$ $\mathcal{O}\!\big(\text{batch\_size}\times(\text{in\_dim}+\text{out\_dim})+\text{in\_dim}\times\text{out\_dim}+\text{in\_dim}^{2}\big)$
DANP Total 14,000,000,000 $2\times(\text{forward}+R\times z)$ + weight update $XX^{\top} +$ R update $4\times\text{batch\_size}\times\text{in\_dim}\times\text{out\_dim}+4\times\text{batch\_size}\times\text{in\_dim}^{2}+2\times\text{in\_dim}^{3}$ $\mathcal{O}\!\big(\text{batch\_size}\times(\text{in\_dim}+\text{out\_dim})+\text{in\_dim}\times\text{out\_dim}+\text{in\_dim}^{2}\big)$
Table 6: Approximate FLOPS required for a forward pass through a convolutional kernel of $5\times 5\times 3\times 8$ (H, W, chan_in, chan_out), batch size $1000$, for BP, DBP, and DANP
Algorithm FLOPS Brief description of calculation Approx. Memory Required
BP Total 3,600,000 forward pass + backward pass (BP) + weight update $6\times\text{batch\_size}\times\text{in\_dim}\times\text{out\_dim}$ $\mathcal{O}\!\big(\text{batch\_size}\times(\text{in\_dim}+\text{out\_dim})+\text{in\_dim}\times\text{out\_dim}\big)$
DBP Total 38,193,750 forward $R\times z +$ backward (BP) + decorrelation backward (same as forward) + weight update $XX^{\top} +$ R update $6\times\text{batch\_size}\times\text{in\_dim}\times\text{out\_dim}+4\times\text{batch\_size}\times\text{in\_dim}^{2}+2\times\text{in\_dim}^{3}$ $\mathcal{O}\!\big(\text{batch\_size}\times(\text{in\_dim}+\text{out\_dim})+\text{in\_dim}\times\text{out\_dim}+\text{in\_dim}^{2}\big)$
DANP Total 38,193,750 $2\times(\text{forward}+R\times z)$ + weight update $XX^{\top} +$ R update $4\times\text{batch\_size}\times\text{in\_dim}\times\text{out\_dim}+4\times\text{batch\_size}\times\text{in\_dim}^{2}+2\times\text{in\_dim}^{3}$ $\mathcal{O}\!\big(\text{batch\_size}\times(\text{in\_dim}+\text{out\_dim})+\text{in\_dim}\times\text{out\_dim}+\text{in\_dim}^{2}\big)$
Table 7: Empirical memory use for parameters in MB in a 3-layer fully connected network trained on CIFAR-10
Algorithm Memory Use (MB)
BP 20.04
DBP 68.04
ANP 20.04
DANP 68.04